### v2

# Hierarchical Text Classification  (HTC) vs. eXtreme Multilabel Classification (XML): Two Sides of the Same Medal

Abstract.
Assigning a subset of labels from a fixed pool of labels to a given input text is a text classification problem with many real-world applications, such as in recommender systems. Two separate research streams address this issue. Hierarchical Text Classification (HTC) focuses on datasets with smaller label pools of hundreds of entries, accompanied by a semantic label hierarchy. In contrast, eXtreme Multi-Label Text Classification (XML) considers very large label pools with up to millions of entries, in which the labels are not arranged in any particular manner. However, in XML, a common approach is to construct an artificial hierarchy without any semantic information before or during the training process. Here, we investigate how state-of-the-art models from one domain perform when trained and tested on datasets from the other domain. The HBGL and HGLCR models from the HTC domain are trained and tested on the datasets Wiki10-31K, AmazonCat-13K, and Amazon-670K from the XML domain. On the other side, the XML models CascadeXML and XR-Transformer are trained and tested on the datasets Web of Science, The New York Times Annotated Corpus, and RCV1-V2 from the HTC domain. HTC models, on the other hand, are not equipped to handle the size of XML datasets and achieve poor transfer results. The code and numerous files that are needed to reproduce our results can be obtained from https://github.com/FloHauss/XMC_HTC.
NLP, text classification, hierarchical text classification, extreme multi-labeling
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing Document analysisccs: Computing methodologies Information extraction
1. Introduction
Text classification is an essential task in text mining and natural language processing (NLP), which involves assigning pre-defined labels to input text, with applications ranging from recommender systems and information retrieval to tasks such as sentiment analysis, e-commerce, advertising and news classification (Gasparetto et al., 2022; Chang et al., 2021; Jain et al., 2019; Sun and Lim, 2001a). Since the introduction of text classification decades ago, many approaches have been developed, with a recent trend focusing on deep learning models (Li et al., 2020). Within the landscape of multi-label text classification, two main categories of approaches emerged: Hierarchical Text Classification (HTC) and eXtreme Multi-Label Text Classification (XML). Both approaches can be seen as distinct and relevant methodologies addressing specific issues and challenges in text classification.
HTC is a text classification problem tailored to exploit the inherent information of a hierarchically structured labelset (Zhou et al., 2020a). Furthermore, HTC utilizes a top-to-bottom approach for path construction and hierarchical relationships to improve classification performance (Sun and Lim, 2001b). A common approach to taxonomic hierarchy in HTC is to model a graph or tree, where each node represents a label to be classified (Wang et al., 2022a). In the field of HTC, there have been many different approaches that incorporate hierarchical information. For instance, independently processing text and structure of label hierarchy and then merging the two elements in a joint representation (Chen et al., 2021a). By doing so, it is aimed to reach a combined representation of text and hierarchy that is beneficial for classification (Zhou et al., 2020a). Another usage of hierarchy for text classification within HTC can be narrowed down to local and global approaches  (Zhou et al., 2020b), where local approaches aim to create classifiers at each level of the hierarchy by using strategies such as CNNs. On the other hand, global strategies approach HTC as a form of multi-label classification by integrating hierarchical information directly into models to rely on the inherent structure and relationships within the hierarchy to improve classification accuracy (Sun and Lim, 2001a).
Unlike HTC, XML is designed to handle large datasets and to work with a large number of labels. XML is an advanced text classification domain that focuses on assigning a relevant subset of instances from a large labelset, where the number of labels poses serious computational and statistical challenges, making it difficult to create an XML method that is efficient to train and accurate (Zhang et al., 2024; You et al., 2019). XML methods can be broadly categorized into three areas: sparse linear models, embedding-based methods, and transformer-based models (Zhang et al., 2024). Initially, sparse linear models laid the foundation. They use linear algorithms and efficient partitioning to work with extensive label spaces by implementing hierarchical trees and multi-label classification techniques such as one-versus-all (Wydmuch et al., 2018). Subsequently, embedding-based techniques were introduced to work with neural networks and aimed to learn semantic representations, often combined with sampling for scalability (You et al., 2019). The latest development, transformer-based models, exploits the capabilities of pre-trained transformers to capture complex semantic relationships within the data (Zhang et al., 2024).
Motivation
So far, it is unclear how the XML and HTC techniques would perform in unconventional settings, namely when XML methods encounter hierarchical data and HTC methods work with vast amounts of unstructured labels. Driven by a cross-domain evaluation, this study aims to identify potential limitations and opportunities for improvement within HTC and XML approaches and to assess ways in which the methodologies can be extended to accommodate more diverse datasets. The investigation is expected to provide key information regarding the performance and usability of HTC and XML approaches and potential new applications in the field of multi-label text classification.
Problem Statement
Until now, both HTC and XML have co-existed as two distinct domains of research in the world of multi-label text classification. Both approaches aim to train a classifier on a given dataset containing texts and their respective class labels to predict relevant labels for an input. However, what distinguishes these two approaches is, firstly, that HTC models are additionally provided with a fixed label taxonomy. XML methods create such a hierarchy through clustering at runtime. Secondly, the number of labels in the HTC datasets is much smaller than the number of labels in the XML datasets. A few hundred in the case of HTC datasets, but up to millions when it comes to XML. For that reason, there has been no research or knowledge about a model performance, when applied to datasets from the other domain. To gain a deeper understanding of the individual capabilities of state-of-the-art models from both sides, this separation has to be dissolved by making the datasets from one side accessible to models from the other, without structural information loss or unfeasible spatial or computational costs. A further distinction is that the evaluation measures differ between the HTC domain and the HTC domain, HTC commonly employs standard classification metrics such as F1-micro and F1-macro, whereas XML uses ranking-based metrics such as P@k.
Contribution
We have developed a novel approach that allows cross-domain comparison of HTC and XML models with datasets traditionally used for text classification tasks from the other side. This signifies that we have transferred HTC models to be applicable to XML datasets and vice versa. Furthermore, the R-Precision metric has been applied to all models to ensure a fair comparison. Our results show that XML models, specifically CascadeXML, achieve comparable or better results than HTC models when tested on datasets from both domains for the performance measures P@k and R-Precision. This allows us to argue that XML models are more versatile. The results of our study suggest that models need to be tested more frequently in cross-domain tasks and that at least XML models need to be considered as baselines in HTC research.
Organisation
In Section 2 we summarize related work, subdivided into HTC and XML. In Section 3 we describe how we bring the two worlds together. Here we focus on the modification of benchmark datasets from one world in order to use them in the other. The experimental apparatus is described in Section 4, where preprocessing steps, the general procedure, hyperparameter optimization, and performance measures are explained. An overview of the obtained results is given in Section 5, which are then discussed in Section 6.
2. Related Work
For both, the single and multi-label text classification tasks, fine-tuning pre-trained transformer models yield state-of-the-art performance (Li et al., 2020). They outperform generative large language models such as GPT-4 (OpenAI, 2023) and Llama-2 (Touvron et al., 2023) in accuracy while offering advantages of smaller model size and less training time (Sun et al., 2023; Yu et al., 2023). They also improve on recent graph-based methods (Galke et al., 2022). Therefore, we focus on models based on pre-trained transformers as well as common baseline methods. Below, we discuss the state-of-the-art in both domains separately, first in HTC, followed by XML. For further reading, we refer to surveys like Zangari et al. (2023).
2.1. Hierarchical Text Classification (HTC)
Given a hierarchy of class labels organized as a tree (in short, a label tree) and an input text, the task of HTC is to produce a set of labels from the label tree that accurately represents the input text. Approaches in this field are often distinguished as local and global (Zhou et al., 2020b), where local approaches focus on building appropriate classifiers for each node or each level in the label tree, whilst global approaches employ a single classifier for the whole hierarchy instead. The models we selected generally follow the global approach.
HGCLR (Wang et al., 2022a) is a recent model that integrates the label hierarchy directly into the text encoder using a contrastive learning framework. It employs BERT (Devlin et al., 2019) as the text encoder, which processes the input token sequence and produces hidden representations for each token, with the [CLS] token summarizing the entire sequence for classification tasks. As a graph encoder, HGCLR employs a modified Graphormer (Ying et al., 2021) to capture hierarchical structures. The modified Graphormer integrates spatial encoding, edge encoding, and centrality encoding into its architecture. The self-attention mechanism of this modified Graphormer is expressed as follows:
${\bm{A}}_{{\bm{G}}_{ij}}=\frac{({\bm{f}}_{i}{\bm{W}}_{\bm{Q}}^{\bm{G}})({\bm{f}}_{j}{\bm{W}}_{\bm{K}}^{\bm{G}})^{\top}}{\sqrt{{\bm{d}}_{h}}}+c_{ij}+{\bm{b}}_{\phi(y_{i},y_{j})}$
[…]
HBGL (Jiang et al., 2022) builds upon the ideas of HGCLR. The key difference is that HBGL removes the auxiliary graph encoder component and uses BERT (Devlin et al., 2019) for both, text encoding and graph feature encoding. The model consists of two separate training stages using the same BERT model. In the first one, the global hierarchy-aware label embedding module, the label name is tokenized and used to inject the global hierarchy of the hierarchical label tree to create initial label embeddings. Within this step, the BERT encoder is frozen. Learning is performed by using a modified version of the BERT’s masked language modeling (MLM) objective and by using a custom attention mask to represent the hierarchy. The masked label prediction task is treated as multi-label classification, where a masked leaf label must predict itself and its masked sibling leaf labels. This approach prevents BERT from confusing masked leaf labels under the same parent label. In the second stage, which is the local hierarchy-aware text encoder, the BERT encoder receives three inputs: The appropriate text tokens, the local hierarchy, and the masked local hierarchy. The local hierarchy consists of one token for each hierarchy level, representing an aggregation of the relevant label embeddings within the respective hierarchical level. Likewise, the masked local hierarchy consists of additional masked tokens, one for each of the hierarchy tokens. The model learns to predict this local hierarchy in an autoregressive fashion. Learning is done in the manner of s2s-ft (Bao et al., 2021), again with a custom attention mask to prevent higher levels in the hierarchy from attending lower levels. The predicted local hierarchy can then be used for inference.
Other models that leverage hierarchical information in the context of text classification and work similarly to HGCLR, such as HPT (Wang et al., 2022b), use different graph encoders, such as GAT (Velickovic et al., 2018), and different learning approaches. For example, they focus more on MLM rather than contrastive learning. Certain methods also do not rely on pre-trained transformer models like BERT and instead utilize transformer mechanisms such as attention, as seen in HiAGM (Zhou et al., 2020b). However, research has demonstrated that leveraging pre-trained transformers generally enhances performance. This stance is supported by findings presented by Galke et al. (2022), and further asserted by earlier studies such as HTCInfoMax (Deng et al., 2021) and HiMatch (Chen et al., 2021b), which have explored various methodologies for integrating textual and hierarchical data. Overall, pre-trained transformer-based models seem to achieve state-of-the-art results in the field of HTC. From the mentioned models we select HBGL and HGCLR for our evaluation as both of them use the same datasets and metrics, and show the best performance.
2.2. eXtreme Multilabel Classification (XML)
XML, like HTC, aims to assign a subset of labels to a given text and leverages a hierarchy to improve accuracy and reduce training complexity. However, there are two main differences: Firstly, the hierarchies are not given as additional input but are created by the model itself based on vector representations of the labelset. Secondly, the XML label spaces are much larger than their HTC counterparts.
There are two avenues in which XML techniques can be grouped based on input: Firstly, there are conventional machine learning methods that employ sparse features such as bag-of-words (BoW) features from text. Secondly, there are deep learning methods utilizing the raw textual data. XML algorithms in the first category can be divided into the following three subclasses: one-vs-all (Babbar and Schölkopf, 2017; Yen et al., 2017), label-tree-based (Prabhu et al., 2018), and embedding-based (Bhatia et al., 2015) methods. Recently, deep learning methods from the second category that operate on the full sequence of text inputs (rather than a bag of words) have become popular. They fine-tune pre-trained transformer models like BERT, RoBERTa (Liu et al., 2019), or XLNet (Yang et al., 2019) and have shown promising results in both prediction accuracy and model size. In the following, we will discuss the architectures and strengths and weaknesses of some recent approaches.
AttentionXML (You et al., 2019) as a label-tree-based deep learning model combines two such techniques to take advantage of both approaches. It does not yet use the transformer architecture, instead a BiLSTM (Schuster and Paliwal, 1997) is used. In addition to that, AttentionXML employs two features: The first is a multi-label attention layer, to incorporate different contextual information from the raw input for each label, alongside the semantic information provided by using raw text instead of bag-of-words features. The second one is the creation and usage of a shallow and wide probabilistic label tree for its training, to be able to handle the enormous label spaces present in XML tasks and thus improve the scalability of datasets with a large pool of labels. Through its usage of said tree and the multi-label attention, AttentionXML also shows good performance in predicting rare labels.
To improve efficiency and accuracy LightXML (Jiang et al., 2021) employs end-to-end training and dynamic negative label sampling via generative cooperative networks, allowing the label ranking part to learn progressively from easy to hard instances, preventing overfitting and improving overall model performance. Unlike static negative sampling methods such as AttentionXML, the dynamic negative sampling approach ensures that negative labels are re-sampled for each training instance. LightXML also makes use of pre-trained transformers to obtain text embeddings. The base models of BERT, RoBERTa, and XLNet are used to generate multi-layer text representations by concatenating the hidden state of the [CLS] token of the last five layers of the encoder, which speeds up convergence. Although LightXML outperforms AttentionXML in terms of model size and training time, it scales poorly compared to newer models described below, especially on large datasets.
To address the issue of scalability and the long training times required for the largest datasets, XR-Transformer (Zhang et al., 2024) introduces a recursive architecture. Instead of selecting a fixed label resolution for training, the pre-trained transformer model is gradually refined by increasing label resolutions. This creates a series of simple to difficult training tasks and reduces training costs. To accomplish this, a Hierarchical Label Tree (HLT) is formed by repeatedly applying a $k$-means clustering algorithm on the label features. The feature of a label $l\in L$ is computed by averaging the TF-IDF (Term Frequency-Inverse Document Frequency) representations of the documents that carry a label $l$. Subsequently, the model is trained on each level of the tree in a top-down manner. This multi-resolution learning approach is also studied in other machine learning contexts, such as computer vision (Zhou and Liang, 2023). At each level, a three-step algorithm is executed. The first step, shortlisting, involves selecting the top-$k$ clusters from the parent level. These clusters are then used in the second step, fine-tuning, where the model is fine-tuned on the label resolution of the current level. Finally, in the third step, ranking, the top-$k$ clusters from the current level are ranked. XR-Transformer has been demonstrated to reduce the training times of LightXML by an average factor of 17 while maintaining similar accuracy results.
CascadeXML (Kharbanda et al., 2022) was introduced to address the complexity of XR-Transformer and the scalability of LightXML. It is an end-to-end trainable XML algorithm that outperforms the aforementioned approaches on the P@k metric. The model uses the same ensemble settings as LightXML and XR-Transformer. It utilizes an ensemble of BERT, RoBERTa, and XLNet for smaller datasets and an ensemble of three BERT models with different seeds for random initialization for larger datasets. CascadeXML takes advantage of the multi-layer architecture of transformers and does not treat them as black-box encoders. Instead, CascadeXML trains a single transformer across multiple resolutions, allowing it to use attention maps specific to the tree resolution. At its core, the method clusters all the given labels and creates an HLT – similar to the XR-Transformer. A shortlisting process is performed to select the highest-scoring metalabels, which are referred to as level-one metalabels. This process is continued until the final level is reached, which, in contrast to XR-Transformer, carries out classification at full label resolution.
Inspired by InceptionNet, InceptionXML (Kharbanda et al., 2023) proposes a lightweight architecture for extreme short text classification. It uses an embedding enhancement module to make representations independent of word order. The architecture employs three parallel branches of 1D convolution layers with different filter sizes, followed by an extreme linear classifier. The model integrates skip connections, non-linear activation, and one-vs-all classifiers for label prediction. Training involves end-to-end optimization using binary cross-entropy loss. Overall, InceptionXML achieves state-of-the-art results in extreme multi-label classification for short text queries.
2.3. Transferability
Although HTC and XML both work with multiple labels, HTC methods require input labels to be structured in a hierarchical tree. Therefore, HTC and XML use completely different benchmark datasets. On the HTC side, widely used benchmark datasets are Web of Science (WoS) (Kowsari et al., 2017), The New York Times Annotated Corpus (NYT) (Shimura et al., 2018), and RCV1-v2 (Lewis et al., 2004). On the other hand, Wiki10-31K (Zubiaga, 2012), AmazonCat-13K (McAuley and Leskovec, 2013), and Amazon-670K (McAuley and Leskovec, 2013) are commonly used in the XML literature. Note that these datasets are different in size and evaluation strategy for HTC and XML models. HTC models use datasets with smaller labelsets and fewer data points than XML datasets and evaluate performance by calculating Micro-F1 and Macro-F1 scores. XML models are evaluated with ranking-based metrics, such as P@k and PSP@k (Jain et al., 2016) metrics. This also poses issues related to the comparability of the two worlds. Micro-F1 and Macro-F1 scores are calculated based on hard yes or no classification decisions. P@k and PSP@k are ranking-based metrics that have a somewhat arbitrary cut-off point after the first $k$ predictions, which hinders computing a reasonable F1 score. In order to conduct fair experiments, we implemented the P@k metric for the HTC models and applied the R-Precision metric (Manning et al., 2009) for models from both domains to ensure that the results can be adequately compared. This is described in detail in Section 4.5.
3. Dataset Transfer XML $\leftrightarrow$ HTC
In this section, we discuss our assumptions and methods for converting datasets between the two worlds of HTC and XML. We describe how the HTC hierarchies are collapsed in order to apply XML methods to these datasets (see Section 3.2), and how the label hierarchy is built for HTC models (see Section 3.3), including the adjustment of model parameters and handling of large datasets that exceed the limits of the encoders.
3.1. Assumptions
The datasets used for our experiments were selected from all the datasets mentioned and used by the different models of HTC and XML for text classification. Thus, it is assumed that the datasets used for our experiment are representative of the domains in which they are typically used. Therefore, we expect our experiment to yield meaningful results.
3.2. Transfer of HTC Datasets for XML Models
This direction is the simpler of the two, therefore the transfer can be described briefly. As the XML methods do not require a predefined label hierarchy, the label hierarchies of the HTC datasets can simply be discarded. After this step, there are two ways to complete the transfer. The first option is to transfer all the labels in the labelset and the second option is to transfer only the labels that appear as leaf nodes in the original hierarchy. These are the most specific labels. Note that depending on the hierarchical dataset, these sets are not necessarily disjoint. Our main experiments focus on the first version. Since HTC models train and predict with internal nodes as well as leaf nodes, we found that this version provides the fairest comparison. Nevertheless, we have included the second version as supplementary material in the appendix.
In addition, CascadeXML and XR-Transformer require model-specific input files that are used during training to create their own hierarchies in the form of an HLT (Hierarchical Label Tree, see Section 2). This creation of an HLT is described in detail in the following Section 3.3. We provide these files in our aforementioned GitHub repository.
3.3. Transfer of XML Datasets for HTC Models
In order for XML datasets to work with the HTC models, a hierarchical label tree as constructed by the XML methods is needed to serve as the hierarchy for the HTC methods. In order to generate the hierarchical label tree, we employ a clustering method similar to CascadeXML. Each label in the set is assigned a numerical value based on the feature aggregation of its positive instances (PIFA) (Kharbanda et al., 2022), which is computed as follows. For a label $l_{i}$ and the set $D_{i}$ of documents labeled with $l_{i}$ in the ground truth, PIFA is defined as:
$\operatorname{PIFA}(l_{i})=\sum_{d\in D_{i}}\frac{\text{TF-IDF}(d)}{\operatorname{length}(d)}.$
Subsequently, a balanced $k$-means algorithm is applied recursively on these numeric label representations a total of $n$ times, until a HLT of depth $n$ is created. The difference between our method to transfer XML datasets to HTC models and the application of CascadeXML to the XML datasets is that Equation 3.3 normalizes the TF-IDF vectors by dividing them by the number of words in the documents, while CascadeXML takes the norm of the resulting embedding. This was done to reduce the influence of text length on clustering. This means that labels are clustered together not only when they are relevant to documents with similar features, but also when the labels have similar relevance frequencies. Note, that by generating a hierarchy in this manner, we introduce new labels to the dataset. To allow a meaningful comparison, we use these new intermediate labels in training but remove them during classification. The true labels from the original dataset are now all leaf nodes within the generated hierarchy, even if a label was previously a parent node higher up in the original hierarchy. The branching factors and depths of these trees that were used in our experiments were selected on the basis of CascadeXML. The depth branching factors are summarized in Table 1.
Table 1. Branching factors for each level L1 to L3 (and L4, respectively) of the hierarchical label tree in XML datasets.
Dataset L1 L2 L3 L4
Wiki10-31K 512 8 7.6
AmazonCat-13K 256 8 6.3
Amazon-670K 1024 8 8 10.2
Both HBGL and HGCLR use this hierarchy as input to their underlying encoder within the model. In the case of HBGL, this encoder is BERT, while in HGCLR it is Graphormer. Each label (node) in the HLT is reduced to a node embedding and serves as an input token for the encoder. The tree structure is preserved by using a custom attention matrix, whereby nodes can only attend to successive nodes.
The encoders have an inherent limit regarding the maximum number of input tokens. Since each label acts as an input token, all of the XML datasets surpass this limit. Furthermore, the quadratic complexity of the attention mechanism worsens the issue with large input sizes. To address this limitation, we segment the hierarchical label tree into smaller sub-trees, each maintaining the properties of an HLT.
We perform this segmentation using depth-first search (DFS), exclusively traversing downward to gather just the leaf nodes and their corresponding preceding nodes. This method ensures that each leaf node in the original tree remains a leaf node in the new sub-tree. Additionally, the number of labels in each new sub-tree is kept within the token limit of BERT of 512 tokens.
By doing so, we might lose context between certain nodes, as they will not be aware of each other within the attention mechanism of the encoder. However, large parts of the general hierarchical structure should still be preserved, as distant nodes generally interact less with each other, e. g., an article about sports is less likely to contain a lot of keywords related to astrophysics.
These encoders provide a refined embedding for each label. Although BERT has a maximum input token limit of 512, which might imply a restriction on the number of labels that can be input at once, this is not an issue. For HGCLR, label embeddings are not directly fed into BERT. For HBGL, the label embeddings are aggregated within each hierarchy level, and since the number of hierarchy levels is typically much lower, the classification process is not restricted by the number of labels. However, compressing each label from a hierarchy level within HBGL into a single token will likely diminish the performance of the model.
4. Experimental Apparatus
In the following, we present datasets, preprocessing, procedure, and performance measures.
4.1. Datasets
In total, we consider six different benchmarking datasets. Three of them are from the world of HTC and three of them are from the world of XML. The HTC datasets are Web of Science - WoS by Kowsari et al. (2017), The New York Times Annotated Corpus - NYT by Shimura et al. (2018) and RCV1-v2 by Lewis et al. (2004). The XML datsets are Wiki10-31K by Zubiaga (2012) and AmazonCat-13K, and Amazon-670K each by McAuley and Leskovec (2013). See Table 2 general datasets statistics. The XML datasets can be obtained from “The Extreme Classification Repository: Multi-Label Datasets & Code” by  Bhatia et al. (2016).
Table 2. Summary statistics of the HTC and XML datasets. Amount of labels and data points within each dataset. Avg($L_{i}$) is the average number of classes per sample.
Dataset #Labels #Train #Val #Test Avg($L_{i}$)
HTC
WoS 141 30,070 7,518 9,397 2,00
NYT 166 23,345 5,834 7,292 7,60
RCV1-v2 103 20,833 2,316 781,265 3,24
XML
Wiki10-31K 30,938 14,146 - 6,616 18,64
AmazonCat-13K 13,330 1,186,239 - 306,782 5,04
Amazon-670K 670,091 490,449 - 153,025 5,45
4.2. Preprocessing
For the preprocessing of the datasets, we follow the procedures described within the specific models HGCLR, HBGL, XR-Transformer, and CascadeXML. Additionally, when preparing HTC datasets for the XML models and vice versa, we refer back to the dataset transfer described in Section 3. In addition, to apply the HTC models on the XML datasets AmazonCat-13K and Amazon-679k, the number of training examples is set to $30k$ and another $5k$ for validation, while the test set is left unchanged. This is needed to allow for a reasonable training time.
4.3. Procedure
Our first step is the reproduction of published results for all models we use. With the original results confirmed, we move on to convert the three HTC datasets (WoS, NYT, RCV1-V2) by removing the hierarchy and vice versa injecting a hierarchy into the three XML datasets (Wiki10-31K, AmazonCat-13K, Amazon-670K) via clustering. With all preparations done, all four models are run on the respective datasets adjusted to their domain. All results are then averaged over five individual runs.
4.4. Hyperparameter Optimization
For the initial reproduction of the results for each model, we used the hyperparameters provided by the authors, as those have already been tuned. When it comes to our experiment of applying these models to new datasets, we search for optimal hyperparameter values to adapt a model to the new dataset. However, we do so in a constrained manner by focusing on a few parameters that we expect to possibly have a noticeable impact since optimization through all possible hyperparameters is very time and resource-consuming with high complexity. Hyperparameters we consider for optimization are batch size, number of training epochs, and learning rate.
4.5. Performance Measures
The performance of the models is measured differently in the domains of HTC and XML. HTC models typically use the F1-Micro and F1-Macro metrics. While those are based on classification, XML models use the ranking-based metrics Precision@k (P@k) and Propensity Scored Precision@k (PSP@k) (Jain et al., 2016). F1-Micro and F1-Macro are different ways of aggregating the F1 score (Goutte and Gaussier, 2005). F1-Macro gives equal weight to each class regardless of frequency and F1-Micro gives more weight to frequent classes. F1-Marco is calculated by averaging the F1 score over every class and F1-Micro is calculated by treating every instance equally and computing the harmonic mean of the global (micro) precision and recall across all classes. P@k measures how well a model accurately identifies relevant elements within its top-$k$ predictions. It quantifies how many top-$k$ predictions are also marked as relevant in the ground truth. P@k is formally defined as
$\operatorname{P@k}=\frac{1}{k}\sum_{i=1}^{k}\text{Relevance}(i)$
where Relevance(i) is an indicator function that is $1$ if the $i$-th item is relevant and 00 otherwise. PSP@k does the same, but has a special emphasis on tail labels and rewards their accurate prediction, as they are often challenging to predict (Jain et al., 2016). Tail labels refer to those labels that occur infrequently across the dataset. These labels are part of the long tail distribution, where a few labels are very common (head labels), and many labels are rare (tail labels) (Yue et al., 2024). This metric is particularly relevant in the world of XML because the labelsets are very large, which means that tail labels occur often. In order to enable a meaningful comparison of the performance of models from the HTC and XML domains, it is necessary to define a common evaluation metric. In this context, we decided to adopt the P@k metric from the XML domain and implement it for the HTC models. Furthermore, we use the R-Precision metric, which is an adaptation of the P@k metric, that does not set a fixed $k$ but rather sets $k$ to the number of relevant labels in the ground truth for every data point that is being evaluated (Manning et al., 2009). Formally, R-Precision is defined as
$\operatorname{R\text{-}Precision}=\frac{1}{r}\sum_{i=1}^{r}\text{Relevance}(i)$
where $r$ is the number of relevant labels in the ground truth for the data point that is evaluated.
It is possible to implement these ranking-based metrics for the classification-based HTC models because they work by assigning each label a probability of being relevant to the current query and then classifying by checking if this probability exceeds a certain threshold. This means that they internally also employ a ranking, which can be utilized to calculate P@k scores and the R-Precision. Another possible way of establishing a common metric would have been to adopt the ranking-based XML models to make classification decisions, which would allow the calculation of F1 scores. To do this, one would need to take the relevance ranking produced by the XML models and set a fixed cut-off point above which labels are classified as relevant. However, setting this cut-off point is not trivial and would need to be optimized separately for each dataset. It should also be noted that it is not useful to use a dynamic cut-off point. For example, taking the approach of the R-Precision metric and setting the cut-off point to the number of relevant labels in the ground truth for each query would mean that precision and recall are equivalent, which would make the calculation of F1 scores obsolete (Manning et al., 2009). Furthermore, we excluded the PSP@k metric, due to the considerably smaller size of the HTC labelsets in comparison to the XML labelsets. This makes the problem with the tail labels less relevant for the HTC methods.
The training time of the different models is not taken into consideration in this paper due to the lack of information provided in the original papers and because we had to vary the hardware configuration during experimentation to accommodate specific model-dataset combinations. However, it can be said that XML models generally train substantially faster than HTC models.
5. Results
We report the performance of HTC and XML models on datasets originating from the HTC world. Subsequently, we report the results on datasets from the XML world. The results are the mean of five individual runs for each model-dataset combination.
5.1. HTC and XML methods on HTC datasets
On the NYT dataset (see Table 3, top) CascadeXML performs best on all metrics except P@5 with an R-precision of $84.61$, outperforming both HTC methods and XR-Transformer. It is followed by HGCLR with a difference in R-precision of $0.39$. HGCLR also performs best for P@5 with $71.74$ and a difference of $0.26$ to CascadeXML. For R-Precision, CascadeXML wins $8.68$ points over HBGL and $2.06$ points over XR-Transformer. XR-Transformer also beats HBGL in every metric, improving on HBGL by $6.62$ points in R-Precision. HBGL** is slightly stronger with an XML-style taxonomy than with the standard taxonomy of the dataset. Compared to its standard version, it gains $1.79$ points in R-precision and $2.05$ points in P@1.
Table 3. Results for the HTC datasets, namely New York Times Annotated Corpus (top), Web of Science (middle), Reuters Corpus Volume 1 (bottom). HBGL** indicates that the semantic hierarchy, as provided by the datasets in HTC, has been replaced by a synthetic hierarchical label tree as it is computed in XML.
NY-Times
Method P@1 P@2 P@3 P@5 R-Prec
HGCLR 95.02 93.16 83.98 71.74 84.22
HBGL 83.81 83.26 75.37 64.87 75.93
HBGL** 85.86 84.86 77.08 65.63 77.72
XR-Transformer 92.56 91.50 82.73 70.85 82.55
CascadeXML 95.93 93.56 83.94 71.48 84.61
Web of Science
Method P@1 P@2 P@3 P@5 R-Prec
HGCLR 91.05 86.44 60.27 37.48 86.44
HBGL 85.97 86.02 57.62 34.57 86.02
HBGL** 87.21 86.12 57.49 34.49 86.12
XR-Transformer 81.39 72.98 54.10 35.00 72.98
CascadeXML 88.99 83.38 58.44 36.40 83.38
RCVI-V2
Method P@1 P@2 P@3 P@5 R-Prec
HGCLR 96.78 93.05 83.40 57.86 90.49
HBGL 91.51 88.96 80.43 54.57 85.19
HBGL** 93.92 90.20 80.56 54.08 85.86
XR-Transformer 97.09 92.39 82.36 56.52 88.54
CascadeXML 97.31 92.54 82.61 57.44 89.01
The results on the WoS dataset (see Table 3, middle) show a different picture. HGCLR performs best across all metrics, outperforming HBGL by $0.42$ points, CascadeXML by $3.06$ points, and XR-Transformer by $13.46$ points in R-Precision, respetively. In contrast to the WoS and RCV1 datasets, HBGL has the second-highest R-precision, beating CascadeXML by $2.64$ points. XR-Transformer is well behind all other methods. Note the special nature of the WoS dataset. Each data point has exactly two associated labels, so the precision values for P@2 and R-precision are the same (see Table 2 for details).
The results for the RCV1-V2 dataset (see Table 3, bottom) are similar to those for the NYT dataset. HGCLR performs best on P@2, P@3, P@5, and R-precision with a small margin over CascadeXML. Regarding P@1, CascadeXML has the upper hand with a difference of $0.53$ points to HGCLR. For the other metrics, the improvements of HGCLR over CascadeXML vary from $0.51$ (P@2) to $1.48$ (R-Precision). The differences between CascadeXML and XR-Transformer are rather small for all metrics ($<1.0$ point) and HBGL is outperformed by both methods. Again, the results of HGCLR and CascadeXML are close together, followed by XR-Transformer and HBGL.
5.2. HTC and XML methods on XML datasets
On the Wiki10-31K dataset (results are shown in Table 4, top) XR-Transformer performs best across all metrics with an R-Precision of $38.93$ and is followed by CascadeXML with a difference of $3.17$ points. XR-Transformer beats HBGL by far with differences up to $35.02$ points (P@1) and $12.10$ points in R-Precision.
Table 4. Results for the XML datasets, namely Wiki10-31K (top), AmazonCat-13K (middle), and Amazon-670K (bottom). For methods marked with † no results could be obtained due to resource limitations.
Wiki10-31K
Method P@1 P@2 P@3 P@5 R-Prec
HGCLR - - - - -
HBGL 53.10 52.79 53.15 52.32 26.83
XR-Transformer 88.12 84.87 80.39 70.74 38.93
CascadeXML 86.80 81.90 76.70 67.04 35.76
AmazonCat-13K
Method P@1 P@2 P@3 P@5 R-Prec
HGCLR - - - - -
HBGL 61.01 62.21 63.10 63.20 54.67
XR-Transformer 96.54 90.89 83.38 67.76 82.02
CascadeXML 96.98 91.66 84.54 68.03 83.13
Amazon-670K
Method P@1 P@2 P@3 P@5 R-Prec
HGCLR - - - - -
HBGL - - - - -
XR-Transformer 49.17 46.12 43.75 39.91 39.91
CascadeXML 47.86 45.10 42.88 39.15 39.66
Results on the AmazonCat-13K dataset (results are shown in Table 4, middle) show a slightly different picture in the comparison of the XML models. CascadeXML outperforms both XR-Transformer and HBGL, reaching an R-Precision of $83.13$. Although the improvement upon XR-Transformer is rather small with $1.11$ points, again a sizeable gap to HBGL is observed with a difference of $28.46$ points in R-Precision. Also, the overall performance for this dataset is high compared to other datasets, reaching values in R-Precision up to $83.13$.
The Amazon-670K dataset (results are shown in Table 4, bottom) with its extensive label space represents a greater challenge than the previous datasets. Unfortunately, we were unable to produce results for HBGL due to resource limitations. Even with the adjustments to the segmentation described in Section 3.3, we could not get around the fact that HTC methods are not designed to deal with extremely large label spaces. Comparing the XML models, XR-Transformer beats CascadeXML across all metrics with an R-Precision of $39.91$, slightly outperforming its counterpart by $0.25$ points.
Due to resource limitations, HBGL fails on the Amazon 670k dataset and no results could be produced for HGCLR, too. The latter is because HGCLR faces several critical challenges that need to be addressed. The main problem is that the model uses numerous $O(n^{2})$ complexity functions in different parts of the code, namely in graph encoding and contrastive learning. Here, the inclusion of the extreme number of labels used in XML datasets is likely to cause memory problems on the GPU. With the help of the authors of HGCLR, the problem was narrowed down to HGCLR’s dense attention functions, which are difficult to optimize for large numbers of labels. For this reason, we were unable to complete experiments with HGCLR on XML datasets.
6. Discussion
Section 6.1 discusses the main findings, followed by Section 6.2, which addresses potential threats to the validity of the results. Section 6.3 examines how the results can be generalized in a broader NLP context.
6.1. Key Scientific Insights
Top Performers
Our main takeaway is that while XML methods are well suited for transfer to the HTC domain and achieve competitive results on HTC datasets, HTC methods are not able to manage the sheer size of labels in the XML datasets and are not suitable for use in the XML domain. XML models and their label-clustering techniques work well on HTC datasets and, due to their inherent method of creating an artificial hierarchy, do not even rely on the taxonomies that come with HTC datasets. On the HTC datasets NYT and RCV1-V2, the result of CascadeXML and HGCLR are very close together in first and second place for the P@1 to P@5 metrics and R-Precision, with CascadeXML slightly better on NYT and HGCLR narrowly winning on RCV1-V2.
The only HTC dataset, on which both HTC methods HBGL and HGCLR remain best on the R-Precision metric is the WoS dataset. This is likely due to the unique anatomy of this dataset. While all the other datasets have variable numbers of relevant labels for each document, with an average number of relevant labels of 3.2 and upwards, the WoS dataset contains exactly two relevant labels for each data point and also comes with a specific label hierarchy. Every document is assigned a single parent label and a single leaf label. This makes it more similar to single-label datasets than to multi-label datasets, and it may not be suitable for the clustering techniques used in XML models. It should also be noted that HGCLR outperforms HBGL on all three datasets from the HTC domain. At first glance, this seems to contradict the results presented by the authors of HBGL, which indicate that HBGL is the stronger model of the two, but can be explained by the different selection of performance metrics in the HBGL paper and this one. HBGL produces better F1 scores and HGCLR is stronger on the P@k metric and R-Precision. This means that HGCLR is better at identifying the top labels, but struggles to correctly identify all the relevant labels, resulting in worse F1 scores. HBGL seems to be the opposite. While its top predictions are not as strong as those of HGCLR or CascadeXML, it is better at correctly identifying relevant labels that are less obvious.
For the datasets selected from the XML domain, the comparison is not close. Neither HBGL nor HGCLR produce competitive results. HGCLR is unable to complete the training process on any of the XML datasets due to its $O(n^{2})$ complexity, where $n$ is the number of possible labels. HBGL faces similar problems and does not produce results on the Amazon-670K dataset, even on the $30k$ subset of training samples. For comparison, the largest HTC dataset regarding label space is NYT with $166$ labels. Even the smallest XML dataset in terms of labels, AmazonCat-13K, already contains $13,330$ labels. While HBGL is able to complete its training and testing pipelines on the smallest datasets from the XML world - AmazonCat-13K and Wiki10-31K - it still performs considerably worse than the models from the XML domain. Between these two, XR-Transformer wins the comparison on the Wiki10-31K and Amazon-670K datasets, and CascadeXML performs best on AmazonCat-13K.
Clustering Techniques
While both the HTC and XML methods belong to the hierarchical multi-label text classification domain, which leverages relational information about the label-space, their respective approaches to utilizing this information are quite different. The hierarchies employed by HTC datasets and methods differ from the HLTs created prior to the training process of XML models. While the former are based on semantic similarities, the latter are characterized by clustering labels together that have similar relevance frequencies and are relevant on documents with similar TF-IDF feature representations. The results of our experiments indicate that the former approach, which is oriented towards a human understanding of label similarity, may not be as beneficial for a machine-driven learning process as the latter. Furthermore, XML models are particularly strong on the P@1 metric, which could also be due to this difference in clustering strategy. Frequency-based clustering encourages the models to predict frequent labels even more often, which is beneficial for predicting head labels and increases P@1, especially when the evaluated dataset contains strong head labels that occur disproportionately, as is the case in the RCV1-V2 and NYT datasets. This is also supported by the results of HBGL** where the native taxonomy was replaced by an XML-style one. Although the improvements made are marginal and too small to claim that this version of clustering is superior in a broader machine-learning context, they are best in P@1.
Metrics
As mentioned in Section 4.5, the XML and HTC domains use different measures to evaluate model performance. In the domain of HTC, F1 scores are mainly used, and in the domain of XML, the P@k metric is predominant. However, the results of our experiments suggest that it is not sufficient to use only one or the other to capture all aspects of model performance. Models should ideally be evaluated on all of these metrics, and possibly even the R-Precision metric.
As presented by the authors of HBGL, HBGL outperforms HGCLR on both the Micro-F1 and Macro-F1 metrics, but in our experiments, HGCLR achieved better results on the P@k and R-Precision metrics, suggesting that HBGL is not stronger per se, but that the two models are rather good at different aspects of text classification. A comparison based on just one set of these metrics would have missed this observation. It would therefore be beneficial to make a comparison using a wider range of metrics, rather than focusing on just one set of them.
It should be noted that some models would need to be adapted to support both families of performance measures, as F1-scores are based on yes or no classifications and both P@k and R-Precision scores are based on rankings. The system under test would have to be able to support both paradigms.
6.2. Threat to Validity
There are a number of threats to validity that need to be discussed. The first relates to the structure of some of the datasets used. In some datasets, texts are labeled with arbitrary labels, that have no significant semantic meaning (i. e., ”!”, ”!!” , ”AE04752”). While the lack of semantic context in these labels may raise concerns about the overall suitability of such datasets for hierarchical classification, it is unlikely a problem for our purposes. Some of the evaluated models, like HBGL, can perform better by incorporating label semantics (using pre-trained embeddings of the textual representation of the label). Ablation studies for HBGL (Jiang et al., 2022) and HGCLR (Wang et al., 2022a) show that the impact of label semantics is relatively small.
Like the model examinations mentioned in Section 2, the datasets used in this study are exclusively in the English language. Consequently, the results presented in this work are solely applicable to the English language domain and may not be directly transferable to other languages.
Detailed hyperparameter optimization, e. g., in the form of a grid search, was out of scope due to the large number of experiments. Instead, we used the hyperparameters tuned by the authors of the respective methods. For the transfer from HTC to XML, we used the hyperparameters of the datasets that came closest to those of the other domain. These were mostly those from Wiki10-31k in the case of XML datasets for HTC methods and those from RCV1-V2 (HBGL configuration) in the case of HTC datasets for XML methods.
Furthermore, the dataset transfer between XML and HTC methods may not be optimal due to differences in their hierarchical structures. XML datasets do not inherently contain hierarchical labels, instead, they generate them through clustering. These clusters include meta labels, which guide the learning process towards the actual labels used for classification - the leaf nodes. HTC methods then apply these generated hierarchies in their models. However, HTC methods expect the entire hierarchy to consist of labels valid for classification. Since only the leaf nodes are actual labels and the rest are meta labels, HTC models are limited to classifying only at the final level of the hierarchy. This restriction may result in diminished effectiveness, as the model cannot classify at each hierarchical level. In other words, the XML datasets are actually only multi-label classification tasks, where the generated hierarchical label tree is a tool and not used to classify the documents on different levels.
Some datasets commonly used in the XML world, such as Wiki-500K and Amazon-3M, were included in the experimental setting in the first place but could not be tested for XR-Transformer and HTC methods due to resource limitations. We ran our experiments on NVIDIA H100 80GB GPUs with 1TB of RAM available but encountered OOM errors when training XR-Transformer on Wiki-500K and Amazon-3M. For the HTC methods, training time was the limiting factor. We were unable to complete the training within the 48-hour window imposed by our computing resource provider.
The performance results reported for CascadeXML by the authors of the model, Kharbanda et al. (2022), on the datasets Wiki10-31k and Amazon-670k could not be achieved. On the Wiki10-31k dataset, our models exhibit a mean deficit of $3.18$ points across the P@1, P@3, and P@5 metrics. On the Amazon-670k dataset, the analogous mean deficit is $3.76$ points. This might be due to the fact, that in contrast to our experiments, CascadeXML was run in an ensemble setting by the authors of the original paper, which can increase performance. On AmazonCat-13K, however, our numbers of CascadeXML are actually slightly better than the ensemble setting. The influence of the choice of language model(s) used and the impact of using ensemble should be analyzed in future work. The principal outcome of our analysis comparing HTC versus XML methods on the respective datasets is not affected because the CascadeXML used here without ensemble outperforms XR-Transformer on all HTC datasets and the AmazonCat-13k dataset.
6.3. Generalization
Our evaluation results on HTC datasets show that XML methods, in particular CascadeXML, are very competitive and can be considered for hierarchical text classification tasks. CascadeXML performed particularly well on the NYT dataset, even outperforming state-of-the-art HTC methods such as HGCLR and HBGL on some datasets. This suggests that the structure of the NYT dataset, which involves complex and deep hierarchical categorizations of news articles, is particularly well suited to the strengths of XML methods. At the same time, despite the challenging comparison, HGCLR has shown that HTC methods remain relevant and effective in certain scenarios. The results suggest that HTC methods could benefit from adaptation to larger datasets, which would require model refinements to cope with increasing complexity.
7. Conclusion
We investigated the cross-domain capabilities of Hierarchical Text Classification (HTC) and eXtreme Multi-Label Classification (XML) models, using cross-dataset transfer to evaluate their performance on unfamiliar datasets and label spaces. Our results show that XML models, in particular CascadeXML, demonstrate considerable adaptability and competitive accuracy when applied to HTC datasets, effectively handling smaller hierarchical structures despite being designed for large, unstructured label pools. However, the HTC models struggled with the sizes and sparsity of the XML datasets, suggesting scaling limitations.
Our results call for a rethinking of common clustering techniques for XML and HTC methods. Label clustering based on the features of the documents to which they are relevant could potentially be replaced or combined with a label frequency-based clustering approach. Other clustering techniques, such as neural clustering, could also be explored. Lastly, building on the initial success of XML models in the HTC world, extending the application of XML methods to other domains could be beneficial.
Acknowledgements.
The authors acknowledge support from the state of Baden-Württemberg through bwHPC.
Appendix
Appendix A Level-wise Number of Labels in the HTC Datasets
Table 5 shows the details of the hierarchical structure of the HTC datasets.
Table 5. Absolute number of labels in each hierarchy layer of the HTC datasets. Notice that only HTC datasets provide an inherent hierarchical label structure, while XML methods compute the label tree via clustering.
Dataset L1 L2 L3 L4 L5 L6 L7 L8
WoS 7 134
NYT 4 27 51 47 17 12 6 2
RCV1-v2 4 55 43 1
Appendix B Results for Leaf Runs
The transfer of HTC datasets to the XML world can be achieved in a number of ways. In the HTC world, datasets come with semantic label hierarchies, and if a label at the leaf level of that hierarchy is relevant to a query, this means that all parent labels of that leaf label are implicitly relevant to that query as well. In our main experiments, we included both leaf labels and all implicit labels in the transfer of HTC datasets to the XML domain. In this additional analysis, we explore the other possible way of transferring the dataset, which ignores the implied positive labels derived from the hierarchy and focuses only on those labels that have no further children. To enable a fair comparison between XML and HTC models, the datasets were also adjusted for the HTC world, including removing all positive labels that are not leaf labels for each data point. It should be noted that this adjusted task is much harder for models from both worlds for two reasons. First, the number of relevant labels for each query is lower in this version, which has the consequence that P@k results are naturally worse. Second, information is lost in the discarding of labels that are not leaf labels in the hierarchy. The results of these runs with modified datasets can be seen in the following tables. On the NYT dataset, CascadeXML and HGCLR remain the best-performing models, with HBGL’s performance dropping strongly from $75.93$ to $42.30$ points in R-Precision compared to the main analysis. On the Web of Science dataset, which is a single-label dataset after the removal of all parent labels, HBGL overtakes HGCLR as the best-performing model with an R-Precision of $80.65$, narrowly beating out CascadeXML with an R-Precision of $77.47$. XR-Transformer achieves $69.97$ points and HGCLR is in last place with only $22.74$ points. On the RCV1-V2 dataset, HGCLR performs best again followed by CascadeXML and XR-Transformer. HBGL is not able to handle this version of the RCV1 dataset well and achieves an R-Precision of only $8.94$ points. Overall the results of HTC models on these adjusted datasets are very inconsistent. This is likely due to their reliance on relational information between labels. Part of this information is lost when only taking leaf labels into account in training and testing. XML models do not seem to rely on this information as heavily.
Table 6. New York Times Annotated Corpus
NY-Times
Method P@1 P@2 P@3 P@5 R-Prec
HBGL 57.46 42.73 33.10 21.97 42.30
HGCLR 82.93 64.74 53.51 38.82 74.09
XR-Transformer 82.14 62.95 51.78 37.67 72.27
CascadeXML 82.89 63.75 51.99 37.90 74.32
Table 7. Web of Science
Web of Science
Method P@1 P@2 P@3 P@5 R-Prec
HBGL 80.65 40.49 26.99 16.19 80.65
HGCLR 22.74 20.08 18.93 19.74 22.74
XR-Transformer 69.97 40.53 28.53 17.86 69.97
CascadeXML 77.41 42.14 29.03 17.92 77.47
Table 8. Reuters Corpus Volume 1
RCV1-V2
Method P@1 P@2 P@3 P@5 R-Prec
HBGL 13.02 7.50 5.17 3.12 8.94
HGCLR 92.70 59.99 43.81 27.96 83.26
XR-Transformer 80.97 52.56 38.81 25.15 76.38
CascadeXML 86.36 55.39 40.55 26.16 81.75
Appendix C Additional Results for Wiki10-31K and Amazon-670K
Table 9 shows additional results for the Wiki10-31K dataset regarding P@10 and P@20, since the dataset has a higher average number of labels for sample. The general trend of the P@k metric continues also for higher values of $K$. Table 10 shows additionally the P@10 scores for Amazon-670K.
Table 9. Additional results for Wiki10-31K. For methods marked with † no results could be obtained due to resource limitations.
Wiki10-31K
Method P@1 P@2 P@3 P@5 P@10 P@20 R-Prec
HGCLR - - - - - - -
HBGL 53.10 52.79 53.15 52.32 42.95 27.71 26.83
XR-Transformer 88.12 84.87 80.39 70.74 53.82 36.76 38.93
CascadeXML 86.80 81.90 76.70 67.04 50.07 33.67 35.76
Table 10. Additional results for Amazon-670K. For methods marked with † no results could be obtained due to resource limitations.
Amazon-670K
Method P@1 P@2 P@3 P@5 P@10 R-Prec
HGCLR - - - - - -
HBGL - - - - - -
XR-Transformer 49.17 46.12 43.75 39.91 26.83 39.91
CascadeXML 47.86 45.10 42.88 39.15 26.19 39.66